# Test-Time Meta-Adaptation with Self-Synthesis

### 03 Mar 2026 v1

@@ 3.2 Results
  • Base Test-Time Self-Synthesis (TT-SS): The base model generates $k$ synthetic training examples, performs a LoRA update, then attempts the target task.
@@
In the evaluations in Table 1, MASS achieves the strongest performance, substantially outperforming all baselines and improving on the Base model by 15.4pp (x1.35). Without meta-learning, the model shows limited capabilities in generating targeted beneficial synthetic data for self-adaptation (Base-TT-SS), but is able to improve from Base by 3.0pp. Naive TTT slightly hurts performance, suggesting that generic test-time updates without problem-specific supervision can introduce drift. MASS performs strongly in both settings where it updates according to golden solutions and verified self-generated solutions.

### 08 Mar 2026 v2

@@ 3.2 Results
  • Base Test-Time Self-Synthesis ($T$-$S$): The base model generates $k$ synthetic training examples, performs a LoRA update, then attempts the target task.
@@
In the evaluations in Table 1, MASS achieves the strongest performance, substantially outperforming all baselines and improving on the Base model by 15.4pp (x1.35). Without meta-learning, the model shows limited capabilities in generating targeted beneficial synthetic data for self-adaptation (Base-$T$-$S$), but is able to improve from Base by 3.0pp. Naive TTT slightly hurts performance, suggesting that generic test-time updates without problem-specific supervision can introduce drift. MASS performs strongly in both settings where it updates according to golden solutions and verified self-generated solutions.